OpenStudy (anonymous):

For what values of the constant c is the function continuous on (-infinity, +infinity)?

OpenStudy (anonymous):

\[f(x)=\left\{ cx+1, if x \le 3 \right\} and \left\{ cx^2-1, if x >3 \right\} \] it is all supposed to be in one system of equations but I don't know how to format that.

myininaya (myininaya):

looking at cx+1 and cx^2-1 by themselves these functions are both continuous everywhere because they are polynomials but to find c such that f is continuous everwhere you have to find when the left limit will equal the right limit as x approaches 3

OpenStudy (anonymous):

Right, so how would I go about doing that. It says for what values so there is more than 1. Im kinda lost.

myininaya (myininaya):

have you found the left limit?

myininaya (myininaya):

we are trying to solve this equation \[\lim_{x \rightarrow 3^-}f(x)=\lim_{x \rightarrow 3^+}f(x)\]

myininaya (myininaya):

what function is to the left of 3?

OpenStudy (anonymous):

cx+1

myininaya (myininaya):

\[\lim_{x \rightarrow 3^-}f(x)=\lim_{x \rightarrow 3^+}f(x) \\ \lim_{x \rightarrow 3^-}(cx+1)=\lim_{x \rightarrow 3^+}(cx^2-1) \\ \]

myininaya (myininaya):

and the right function would have to be cx^2-1

myininaya (myininaya):

now compute the limit that is on the left hand side of the equation

OpenStudy (anonymous):

ok so it would be 3x+1? or really close to 3?

OpenStudy (anonymous):

3c*

myininaya (myininaya):

why do you replace c with 3 when x is the one that approaches 3

myininaya (myininaya):

ok yes 3c+1

myininaya (myininaya):

so the other side would be ?

OpenStudy (anonymous):

9c-1

myininaya (myininaya):

\[\lim_{x \rightarrow 3^-}f(x)=\lim_{x \rightarrow 3^+}f(x) \\ \lim_{x \rightarrow 3^-}(cx+1)=\lim_{x \rightarrow 3^+}(cx^2-1) \\ 3c+1=9c-1 \]

myininaya (myininaya):

so we need to solve this linear equation

myininaya (myininaya):

so there will be only one value of c

OpenStudy (anonymous):

OH MY GOSH THANKS THAT MAKES SO MUCH SENSE. I DIDN"T REALLY UNDERSTAND WHAT IT WANTED ME TO DO.

myininaya (myininaya):

np